OpenStudy (anonymous):

Solve 2x2 + 12x = -10 by completing the square

OpenStudy (anonymous):

You mean \[2x^{2} + 12x = -10\] ?

OpenStudy (anonymous):

Then \[x ^{2} + 6x= -5\] \[x ^{2} + 6x + 9 = 4\] \[(x + 3)(x + 3) = 4\] \[(x + 3)^{2} = 4\] \[x + 3 = 2\]

OpenStudy (anonymous):

x = -1

OpenStudy (anonymous):

yes that's what I meant! thanks so much(: